# Deep Backtracking Counterfactuals for Causally Compliant Explanations

### v3

@@ 4 Experiments
In our experiments, we contrast DeepBC with existing ideas and showcase the properties and abilities outlined in Section § 3.5. As showing a single example per counterfactual is more illustrative than sampling multiple examples, we mainly focus on mode DeepBC (§ 3.1.2, § 3.3.2) and refer to this variant when using the term DeepBC from this point on. Technical details about the implementation are provided in App. B.
@@ 4.1 Morpho-MNIST
Figure 4: Comparison of backtracking and interventional counterfactuals on MorphoMNIST. The blue shaded areas indicate the probability density of the data in latent (a) or observed (b), (c) space. (a) Given a factual realization (red dot), varying the values of the antecedent $i^{*}$ changes both $u^{*}_{I}$ and the upstream variable $u^{*}_{T}$. Since interventional counterfactuals do not perturb the latents, only the backtracking solution (grey dots) is shown. (b) Interventional counterfactuals (green dots), in contrast to backtracking counterfactuals, leave $t^{*}$ unchanged when the effect variable intensity is taken as antecendent. (c) When treating thickness as the antecedent, counterfactual and backtracking counterfactuals yield identical solutions.
@@
We use Morpho-MNIST, a modified version of MNIST proposed by Castro et al. (2019), to showcase how deep backtracking contrasts with its interventional counterpart (Pawlowski et al., 2020). The data set consists of three variables, two continuous scalars and an MNIST image of a handwritten digit, all depicted in Fig. 1. The first scalar variable $T$ describes the thickness and the second variable $I$ describes the intensity of the digit. They have a non-linear relationship and are positively correlated, as can be seen in Fig. 4 (b) and (c), where the observational density of thickness and intensity is shown in blue. The known causal relationship between thickness and intensity is depicted in Fig. 4 (left); the true structural equations are listed in App. C. We train a normalizing flow for thickness and one for intensity conditionally on thickness, and model the image given $T$ and $I$ via a conditional $\beta$-VAE (Higgins et al., 2017). We here use $d_{i}(u^{\prime}_{i},u_{i})=w_{i}\left\lVert u^{\prime}_{i}-u_{i}\right\rVert_{2}^{2}$ with $w_{i}=1,\;\forall i$ as the default distance function.
@@ Results.
Our results in Figs. 4 and 5 illustrate distinctive properties of the backtracking approach. When choosing the effect variable intensity as the antecedent, backtracking preserves the causal laws and thus changes the upstream (cause) variable thickness accordingly to match the change in intensity as shown in Fig. 4 (a) and (b). This leads to counterfactuals that resemble images from the original data set, where thickness and intensity change simultaneously, as shown in the top row of Fig. 5. DeepBC arrives at these counterfactuals, since $i^{*}\neq i$ can either be achieved by choosing a different $u_{I}^{*}\neq u_{I}$ or by changing the upstream latent $u_{T}^{*}\neq u_{T}$ (because $i^{*}$ also depends on the realization $t^{*}$, which, in turn, depends on $u_{T}^{*}$). As to minimize the sum of squared latent perturbations $d_{T}(u_{T},u_{T}^{*})+d_{I}(u_{I},u_{I}^{*})$, DeepBC tends to change both latent variables from their factual realizations, as can be seen in Fig. 4 (a).
@@
In contrast, the interventional approach breaks the causal link from thickness to intensity when intensity is the antecedent and thus always leaves thickness unchanged, see the green dots in Fig. 4 (b). This results in counterfactual images with high-intensity but atypically low thickness, or low intensity but typically high thickness, as shown in the second row of Fig. 5. In terms of generating counterfactuals that yield faithful insights into the causal relationships underlying the data, this can be considered a weakness of the interventional approach.
@@
However, interventional and deep backtracking counterfactuals can also be identical, as shown in Fig. 4 $\mathbf{(c)}$, where the thickness variable $T$ is used as antecedent. If the antecedent is a root node of the causal graph $G$, as is the case for $T$, the change in $t^{*}\neq t$ cannot be traced back to any latent variable other than $u_{T}$, which is why both $u_{I}^{*}=u_{I}$ and $u^{*}_{\text{Img}}=u_{\text{Img}}$, analogously to interventional counterfactuals. The change in the value $i^{*}$ as a function of $t^{*}$ then solely corresponds to the causal effect of $t^{*}$, for both counterfactuals (Fig. 4 (c)).
As presented in Section § 3, DeepBC further supports sampling via stochastic DeepBC (§ 3.1.1, § 3.3.1) and weighting (§ 3.3.2). In § D.1, we demonstrate these two properties for Morpho-MNIST in Fig. 10 and Fig. 11, respectively.
@@ 4.2 CelebA
We also investigate generating counterfactual celebrity images on the CelebA data set (Liu et al., 2015). The images have a resolution of $128\times 128$ and are annotated with binary attributes {Age, Gender, Beard, Bald}. We adopt the causal graph assumed by Yang et al. (2021b), which is shown in Fig. 6 (top right). Since our optimization algorithms assume differentiability of $\mathbf{F}$ in $\mathbf{u}$ (§ 3.3), we preprocess the data to use the standardized logits of classifiers that we trained to predict each attribute from the image (§ 3.4). Analogously to § 4.1, we then train a conditional normalizing flow for each attribute given its parents in the causal graph, and use a conditional $\beta$-VAE for the image given all four attributes.
@@
In addition to comparing DeepBC with the interventional approach, we also consider the following baselines and ablations in generating counterfactual images:
@@
    Tabular non-causal explanation: Prior work in the field of counterfactual explanations have measured distance directly in terms of $\mathbf{x}$, typically for tabular data such as the low-dimensional attributes of the CelebA data set (e.g., Mothilal et al., 2020; Lang et al., 2022). In the style of these (non-causal) methods (see 2), we train a new regressor that predicts an attribute from all other attributes (not including the image). We do so for each attribute separately. We then employ DeepBC on this regressor, but measure distance in the attributes of $\mathbf{x}$, rather than $\mathbf{u}$.
@@
  • 3)
    Deep non-causal explanation: According to 8, we use an image regressor ($f_{\hat{Y}}$), together with an unconditional image auto-encoder ($f_{X}$) to generate counterfactual explanations. We do so for each attribute ($\hat{Y}$) separately. This corresponds in style to how Jacob et al. (2022); Rodríguez et al. (2021) obtain counterfactual explanations for images. This means that $f_{X}$ does not correspond to the reduced form of a causal model, and so $u_{X}=f^{-1}_{X}(x)$ corresponds to an unstructured embedding space for image data. In contrast to baseline 1) and DeepBC, this approach cannot readily be restricted to only manipulate pre-specified high-level features such as the four attributes of the CelebA data set, without falling back to measuring distance in these variables directly (see baseline 1). For our experiments, we apply DeepBC in this unstructured latent space and use classifiers to extract the high-level attributes from the counterfactual image.
Quantitative Evaluation Metrics.
[…]
The plausible metric measures the probability density of the generated counterfactual attributes, under the data distribution. In this sense, we can think of counterfactuals with high probability density (or low -log density) as being more plausible, i.e., being closer to the data manifold (see Karimi et al. (2022)). The obs metric measures the distance among attributes between factual realization and counterfactual, thereby evaluating closeness in the observation space (regardless of whether the causal laws are preserved). This is the metric that has been optimized in the original work on counterfactual explanations (Wachter et al., 2017). The causal metric measures the difference in exogenous variables between realization and counterfactual, which can be interpreted as the degree of preservation of the causal mechanisms: If changes in latent variables exceed what is minimally necessary under the assumption of retained mechanisms, it suggests that the counterfactual does not effectively preserve the causal mechanisms. We elaborate in greater detail on the reasoning and detailed implementation of the metrics in § C.1.
Table 1: Quantitative Evaluations for CelebA. Three considered metrics $\pm$ standard deviation of the different baselines for $500$ iterations. Lower values are better, best per metric is highlighted in bold. The two distance functions assessed ($m$) are SQU for squared distance, and ABS for absolute distance. More details can be found in § C.1.
Metric m Tabular non-causal Interventional Deep non-causal Wrong graph DeepBC
@@
In § D.2, we further demonstrate the property of DeepBC to take into account antecedents that consist of more than a single variable.
Quantitative Results.
We quantitatively assess multiple properties of sparse mode DeepBC in comparison to the baselines in Tab. 1. While the deep non-causal method achieves the best result on the plausible score, the interventional approach preserves the attribute values the best (obs score). This is likely due to the fact that counterfactuals on leaf nodes of the causal graph in Fig. 6 do not change the values of their parents at all, in contrast to the other methods. Finally, as expected, DeepBC obtains the best score on the causal metric, independent of distance function $m$. This can be explained due to the fact that DeepBC always respects downstream effects (in comparison to other methods), as exemplified in Fig. 6.
@@ 5 Related Work
The most similar existing work to ours is that of Mahajan et al. (2019). Similarly to the present work, the authors also employ deep generative models and measure the distance between factual and counterfactual examples in a latent space. The most distinctive difference to our work is that Mahajan et al. (2019) impose causal constraints via a causal proximity loss in the visible variables $\mathbf{x}$ that assumes additive Gaussian noise, thereby restricting the types of causal mechanisms $f_{i}$ under consideration. The causal proximity loss approach is at odds with the backtracking philosophy (Lewis, 1979) that we follow. In our approach, all changes are traced back solely to latent variables $\mathbf{u}$ that are embedded into the deep causal model, such that all causal constraints are satisfied automatically by construction (§ 3.5). This obviates the need for an additional loss and straightforwardly allows for general noise dependencies (see  App. C). At the same time, our approach is more versatile as any subset of the given variables could be used as antecedent, whereas the method of Mahajan et al. (2019) only supports a specific label variable (see 2), similar to the vast majority of counterfactual explanation methods.
@@ 6 Discussion
In general, neither the structural equations nor the reduced form (see § 2.1) are identifiable from observational data (Karimi et al., 2020; Hyvärinen et al., 2023; Locatello et al., 2019). However, under certain conditions, it has been shown that the structural equations (and therefore the reduced form) can be identified, up to simple transformations (Javaloy et al., 2023). If the causal graph and/or underlying variables are not known either (i.e., we only have access to the high-dimensional images), the problem becomes even more challenging because both the graph and underlying attributes (e.g., the variables beard, gender, etc. in § 4.2) can also not be identified, in general. Recent works have established numerous conditions under which identifiability of variables and/or graph holds up to certain transformations (e.g., Buchholz et al., 2023; Lippe et al., 2023; Liang et al., 2023; von Kügelgen et al., 2023a). We view solutions and results to identifying the causal graph and/or the reduced form as complementary to our proposed method and note that the learned mechanisms of DeepBC generally only correspond to approximations of the true mechanisms.
@@
A possible future line of research could be to explore how backtracking could be implemented for generative models whose latent variables cannot be inferred deterministically from the factual realization, such as diffusion probabilistic models (Ho et al., 2020) and generative adversarial networks (Goodfellow et al., 2014), both of which are not invertible in general. One conceivable solution might be to adapt 4/6 as to jointly sample/optimize over $\mathbf{u}$ and $\mathbf{u}^{\prime}$.
Counterfactuals without Access to the Reduced Form.
The explicit access to the reduced form of a causal model allows for obtaining causally compliant solutions for varying choices of distance functions (§ 3.5, § 4). We note, however, that plenty of non-causal counterfactual methods exist that do not rely on knowledge of the reduced form (see e.g., Guidotti (2022)). In addition, recent work has shown that interventional counterfactuals can be identified for invertible SCMs (Nasr-Esfahany et al., 2023) and it could be that similar results hold for backtracking as well, likely necessitating further restrictions on the choice of backtracking conditional and/or assumptions on the reduced form. This may render the generation of backtracking conditional feasible without access to the reduced form expression, in specific settings.
@@ 7 Conclusion
Our anonymized source code is available at https://anonymous.4open.science/r/DeepBC_review. Detailed instructions for reproducing all experiments are provided in the README.md file at the top level of the repository. All parameters can be found in the config folders within the respective subfolders. In addition, we provide a detailed description of the optimization parameters in § B.1, training procedures in § B.3.1 and deep learning architectures in § B.3.2.

### v4

@@ 4 Experiments
In our experiments, we contrast DeepBC with existing ideas and showcase the properties and abilities outlined in Section § 3.5. As showing a single example per counterfactual is more illustrative than sampling multiple examples, we mainly focus on mode DeepBC (§ 3.1.2, § 3.3.2) and refer to this variant when using the term DeepBC from this point on. When referring to stochastic DeepBC, we always explicitly write stochastic DeepBC. Technical details about the implementation are provided in App. B.
@@ 4.1 Morpho-MNIST
Figure 4: Counterfactual Scalar Variables on MorphoMNIST. The blue shaded areas indicate the probability density of the data. (a) (i) Given a factual realization (red dot), varying the values of the antecedent $i^{*}$ changes both $u^{*}_{I}$ and the upstream variable $u^{*}_{T}$. Since interventional counterfactuals do not perturb the latents, only the backtracking solution (grey dots) is shown. (ii) Interventional counterfactuals (green dots), in contrast to backtracking counterfactuals, leave $t^{*}$ unchanged when the effect variable intensity is taken as antecendent. (iii) When treating thickness as the antecedent, counterfactual and backtracking counterfactuals yield identical solutions. (b) For the correct graph, DeepBC counterfactuals for antecedent thickness do not change as the backtracking conditional (corresponding distance function shown under each subplot) is changed. When we performing DeepBC with the wrong graph ($I\rightarrow T$), causal compliance as described in § 2.3 is violated.
@@
We use Morpho-MNIST, a modified version of MNIST proposed by Castro et al. (2019), to showcase how deep backtracking contrasts with its interventional counterpart (Pawlowski et al., 2020) and how the generated results depend on the correct causal graph (§ 2.1). The data set consists of three variables, two continuous scalars and an MNIST image of a handwritten digit, which all correspond to the observable variables (see § 2.1), depicted in Fig. 1. The first scalar variable $T$ describes the thickness and the second variable $I$ describes the intensity of the digit. They have a non-linear relationship and are positively correlated, as can be seen in Fig. 4 (a) (ii), where the observational density of thickness and intensity is shown in blue. The known causal relationship between thickness and intensity is depicted in Fig. 4 (top left); the true structural equations are listed in App. C. We train a normalizing flow for thickness and one for intensity conditionally on thickness, and model the image given $T$ and $I$ via a conditional $\beta$-VAE (Higgins et al., 2017). We here use $d_{i}(u^{\prime}_{i},u_{i})=w_{i}\left\lVert u^{\prime}_{i}-u_{i}\right\rVert_{2}^{2}$ with $w_{i}=1,\;\forall i$ as the default distance function and note that the $u_{i}$ correspond to $u_{T}$, $u_{I}$ and $u_{\text{img}}$.
@@ Results.
Our results in Fig. 4 (a) and Fig. 5 illustrate distinctive properties of the backtracking approach, in comparison to interventional counterfactuals. When choosing the effect variable intensity as the antecedent, backtracking preserves the causal laws and thus changes the upstream (cause) variable thickness accordingly to match the change in intensity as shown in Fig. 4 (ii) and Fig. 5. This leads to counterfactuals that resemble images from the original data set, where thickness and intensity change simultaneously, as shown in the top row of Fig. 5. DeepBC arrives at these counterfactuals, since $i^{*}\neq i$ can either be achieved by choosing a different $u_{I}^{*}\neq u_{I}$ or by changing the upstream latent $u_{T}^{*}\neq u_{T}$ (because $i^{*}$ also depends on the realization $t^{*}$, which, in turn, depends on $u_{T}^{*}$). As to minimize the sum of squared latent perturbations $d_{T}(u_{T},u_{T}^{*})+d_{I}(u_{I},u_{I}^{*})$, DeepBC tends to change both latent variables from their factual realizations, as can be seen in Fig. 4 (i).
@@
In contrast, the interventional approach breaks the causal link from thickness to intensity when intensity is the antecedent and thus always leaves thickness unchanged, see the green dots in Fig. 4 (ii). This results in counterfactual images with high-intensity but atypically low thickness, or low intensity but typically high thickness, as shown in the second row of Fig. 5. In terms of generating counterfactuals that yield faithful insights into the causal relationships underlying the data, this can be considered a weakness of the interventional approach.
@@
However, interventional and backtracking counterfactuals can also be identical, as shown in Fig. 4 $\mathbf{(iii)}$, where the thickness variable $T$ is used as antecedent. If the antecedent is a root node of the causal graph $G$, as is the case for $T$, the change in $t^{*}\neq t$ cannot be traced back to any latent variable other than $u_{T}$, which is why both $u_{I}^{*}=u_{I}$ and $u^{*}_{\text{Img}}=u_{\text{Img}}$, analogously to interventional counterfactuals. The change in the value $i^{*}$ as a function of $t^{*}$ then solely corresponds to the causal effect of $t^{*}$, for both counterfactuals (Fig. 4 (iii)).
DeepBC depends on the reduced form that is learned from data 1, which in turn depends on the causal graph that is assumed (§ 2.1). In Fig. 4, counterfactuals for the true graph are compared to those from a model that was trained in the same way, with the only difference that the arrow from thickness was reversed. For the correct causal graph ($T\rightarrow I$), we observe that the counterfactuals for antecedent thickness must be invariant with respect to the choice of backtracking conditional (corresponding distance functions are shown below each subplot). This is because intensity is downstream of thickness and so (as mentioned in the previous paragraph and described in § 2.3) it must hold that $u^{*}_{I}=u_{I}$, for any choice of backtracking conditional. However, when using the wrong causal graph ($I\rightarrow T$), we see that the solutions are different and causally incompliant (in the sense of § 2.3). This is because $U_{I}$ is causally upstream of $T$ in the wrong graph and thus $u^{*}_{I}$ contributes to the realization of the antecedent $t^{*}$, in ways that depend on the choice of the backtracking conditional.
As presented in Section § 3, DeepBC further supports sampling via stochastic DeepBC (§ 3.1.1, § 3.3.1), which is demonstrated in Fig. 10. We also present additional experimental results for antecedent intensity using mode DeepBC in Fig. 11, using different distance functions and weightings (§ 3.4).
@@ 4.2 CelebA
We also investigate generating counterfactual celebrity images on the CelebA data set (Liu et al., 2015). The images have a resolution of $128\times 128$ and are annotated with binary attributes {Age, Gender, Beard, Bald}. Both the image and the attributes correspond to the observable variables. We adopt the causal graph assumed by Yang et al. (2021b), which is shown in Fig. 6 (top right). We focus on generating counterfactuals that manipulate the considered variables sparsely. Since our optimization algorithms assume differentiability of $\mathbf{F}$ in $\mathbf{u}$ (§ 3.3), we preprocess the data to use the standardized logits of classifiers that we trained to predict each attribute from the image (§ 3.4). Analogously to § 4.1, we then train a conditional normalizing flow for each attribute given its parents in the causal graph, and use a conditional $\beta$-VAE for the image given all four attributes.
@@
In addition to comparing DeepBC with the interventional approach, we also consider the following baselines and ablations in generating sparse counterfactuals:
@@
    Tabular non-causal explanation: Prior work in the field of counterfactual explanations have measured distance directly in terms of $\mathbf{x}$ 4, typically for tabular data such as the low-dimensional attributes of the CelebA data set (e.g., Mothilal et al., 2020; Lang et al., 2023). In the style of these (non-causal) methods (see 4), we train a new regressor that predicts an attribute from all other attributes (not including the image). We do so for each attribute separately. We then employ sparse DeepBC (§ 3.4) on this regressor, but measure distance directly in the attributes of $\mathbf{x}$, instead of measuring distance in $\mathbf{u}$. In order to generate a counterfactual image in Fig. 6, we use the conditional variational auto-encoder that was trained for DeepBC and condition the auto-encoder on the counterfactual attribute values, while keeping the realizations $u_{\text{Img}}$ from the factual example, i.e., $u^{*}_{\text{Img}}=u_{\text{Img}}$.
@@
We note that deep counterfactual explanation methods 10 are not applicable for rendering sparse changes, because they cannot identify the underlying attribute variables that the counterfactuals should be sparse in.
@@ Results.
In § D.2.1, we further demonstrate the property of DeepBC to take into account antecedents that consist of more than a single variable. We also demonstrate quantitative experiments and their results in § D.2.2.
@@ 5 Related Work
The most similar existing work to ours is that of Mahajan et al. (2019). Similarly to the present work, the authors also employ deep generative models and measure the distance between factual and counterfactual examples in a latent space. The most distinctive difference to our work is that Mahajan et al. (2019) impose causal constraints via a causal proximity loss in the observable variables $\mathbf{x}$ that assumes additive Gaussian noise, thereby restricting the types of causal mechanisms $f_{i}$ under consideration. The causal proximity loss approach is at odds with the backtracking philosophy (Lewis, 1979) that we follow. In our approach, all changes are traced back solely to latent variables $\mathbf{u}$ that are embedded into the deep causal model, such that all causal constraints are satisfied automatically by construction (§ 3.5). This obviates the need for an additional loss and straightforwardly allows for general noise dependencies (see  App. C). At the same time, our approach is more versatile as any subset of the given variables could be used as antecedent, whereas the method of Mahajan et al. (2019) only supports a specific label variable (see 4), similar to the vast majority of counterfactual explanation methods.
@@ 6 Discussion & Limitations
In general, neither the structural equations nor the reduced form (see § 2.1) are identifiable from observational data (Hyvärinen et al., 2024; Karimi et al., 2020; Locatello et al., 2019). However, under certain conditions, it has been shown that the structural equations (and therefore the reduced form) can be identified, up to simple transformations (Javaloy et al., 2023; Nasr-Esfahany et al., 2023). If the causal graph and/or underlying variables are not known either (i.e., we only have access to the high-dimensional images), the problem becomes even worse because both the graph and underlying attributes (e.g., the variables beard, gender, etc. in § 4.2) can also not be identified, in general. Recent works have established numerous conditions under which identifiability of variables and/or graph holds up to certain indeterminancies (e.g., Lachapelle et al., 2022; Buchholz et al., 2023; Lippe et al., 2023; Liang et al., 2023; von Kügelgen et al., 2023a). We view solutions and results to identifying the causal graph and/or the reduced form as complementary to our proposed method and note that the learned mechanisms of DeepBC generally only correspond to approximations of the true mechanisms.
@@
A potential avenue for future research could be to explore how backtracking could be implemented for generative models whose latent variables cannot be inferred deterministically from the factual realization, such as diffusion probabilistic models (Sohl-Dickstein et al., 2015; Ho et al., 2020) and generative adversarial networks (Goodfellow et al., 2014), both of which are not invertible in general (although approximately invertible variants like DDIMs (Song et al., 2021) do exist). One conceivable solution might be to adapt 6/8 as to jointly sample/optimize over $\mathbf{u}$ and $\mathbf{u}^{\prime}$ or to employ variational inference techniques.
Model Misspecification.
The dimensionality of the latent variables for variational autoencoders plays a role in the generated counterfactuals. [^6] For normalizing flows, in contrast to variational autoencoders, the latent space must have the same dimensionality as the observed space. Typically, one assumes this dimensionality to be lower than the dimensionality of the data, motivated by the assumption that the data lives on a low-dimensional manifold (Reizinger et al., 2022; Bonheme & Grzes, 2022; Pope et al., 2021). Another source of model misspecification that may render DeepBC counterfactuals incorrect stems from the non-invertibility of the true underlying mechanisms. While this invertibility is a fairly common assumption in the literature (e.g., Pawlowski et al. (2020); Javaloy et al. (2023); Nasr-Esfahany et al. (2023); Hoyer et al. (2008)), it does not hold in general [^7] A simple example is the assignment $X\leftarrow U^{2}$. and can also not be ruled out from data alone. A remedy for this restriction may be to extend our method to non-invertible generative models, as outlined above.
Limitations of Non-Causal Counterfactual Methods.
The explicit access to the reduced form of a causal model (or at least a good approximation thereof) allows for obtaining causally compliant solutions for varying choices of distance function (§ 3.5, § 4). We note, however, that many non-causal counterfactual methods exist (corresponding to different variants of 4) that do not rely on knowledge of underlying causal variables and their causal relationships (see e.g., Guidotti (2022)). Based on theoretical results in independent component analysis (Hyvärinen & Pajunen, 1999), we suspect that for certain types of backtracking conditional (e.g., rotationally invariant ones) alongside further assumptions on the generative process (e.g., $x_{\text{img}}$ is a conformal map of all latents $\mathbf{u}$) and dimensionality of $\mathbf{u}$, deep counterfactual explanations 10 on $x_{\text{img}}$ may be able to perform causally compliant backtracking implicitly. In the general setting, however, our empirical results (§ 4) show that this is not the case. This can be made concrete on the Morpho-MNIST example (§ 4.1): Neither the underlying variables intensity and thickness, nor their causal relationship can (in general) be implicitly identified (as discussed in the first paragraph of this section), for example by using a deep counterfactual explanation method applied to the image representation 10. In fact, the experiments in Fig. 4 (b) clearly show that backtracking counterfactuals depend on the true causal graph between thickness and intensity.
@@ 7 Conclusion
Our source code is available at https://github.com/rudolfwilliam/DeepBC. Detailed instructions for reproducing all experiments are provided in the README.md file at the top level of the repository. All parameters can be found in the config folders within the respective subfolders. In addition, we provide a detailed description of the optimization parameters in § B.1, training procedures in § B.3.1 and model architectures in § B.3.2.
Acknowledgements
The authors are grateful for valuable discussions with Valentyn Boreiko, Partha Ghosh, and Luigi Gresele. The authors acknowledge the support provided by the German Research Foundation. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Klaus-Rudolf Kladny.